let f(x)=2x+1 evaluate f as indicated f(5+h)
Answer by stanbon(75887)

let f(x)=2x+1 evaluate f as indicated f(5+h)
-----------
f(5+h) = 2(5+h)+1
= 10+2h+1
=11 + 2h
--------------
Cheers,
Stan H.